OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Ocean Tours. : Age Ticket Price Under 3 years free 3 to 12 years $36 Over 12 years $48

OpenStudy (anonymous):

On a certain day, the company took 92 people on whale watching trips. There were 40 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (92 - x)36 + (92 - 40) ⋅ 48 = C (40 + x)36 + (92 - (40 + x)) ⋅ 48 = C (92 + x)36 + (92 - (40 + x)) ⋅ 48 = C (40 - x)36 + (92 - 40) ⋅ 48 = C

OpenStudy (mathstudent55):

There were 92 people. 40 were children. How many were adults?

OpenStudy (anonymous):

is the lat one and why

OpenStudy (mathstudent55):

Can you answer my question? It's part of the explanation. You are correct, btw.

OpenStudy (mathstudent55):

If there are 92 people total and 40 were children, how many were adults?

OpenStudy (anonymous):

40

OpenStudy (mathstudent55):

No. There are 40 children. There are 92 people altogether. There are 92 - 40 adults.

OpenStudy (mathstudent55):

The number of adults is 92 - 40. Ok?

OpenStudy (anonymous):

thanks are u busy i need help i lot

OpenStudy (texaschic101):

do you understand how @mathstudent55 arrived at the answer ?

OpenStudy (mathstudent55):

Ok. The number of children is 40. The number of adults is 92 - 40. Now we are also told that of the 40 children, x number of those children are under 3. There are 40 children x are under 3. How many children are not under 3?

OpenStudy (texaschic101):

I think he just wanted the answer

OpenStudy (mathstudent55):

@texaschic101 He wrote above he knows the answer is d and he wants an explanation. That's what we are working on now.

OpenStudy (anonymous):

no i got the answer by my own

OpenStudy (mathstudent55):

Ok. Can we continue? There are 40 children. x of them are under 3. How many of the children are not under 3?

OpenStudy (anonymous):

70

OpenStudy (mathstudent55):

If there are only 40 children, how can 70 children be over 3?

OpenStudy (mathstudent55):

Let me ask you this. There are 10 boys and girls in a class. 4 are boys. How many are girls?

OpenStudy (anonymous):

6

OpenStudy (mathstudent55):

Right. How did you get 6? You did 10 - 4 Now do the same with 40 and x There are 40 children x are under 3 How many are not under 3?

OpenStudy (mathstudent55):

There are 10 boys and girls in a class. 4 are boys. How many are girls? Answer: 10 - 4 = 6 There are 10 boys and girls in a class. x are boys. How many are girls? Answer: 10 - x There are 40 children. x are under 3. How many are over 3? Answer: 40 - x Read the above and concentrate on the third question.

OpenStudy (mathstudent55):

Ok. I'll finish it. There are 92 people. 92 - 40 are adults who pay $48 per person 40 - x are children over 3 who pay $36 x are children who pay nothing The cost is: C = 36(40 - x) + 48(92 - 40)

OpenStudy (mathstudent55):

Ok, gtg. Go through this. The problem is explained in detail.